Paul picked 2 2/3 bushels of apples. Karl picked 4 1/3 bushels of apples. Carol picked 3 2/3 bushels of apples. How many bushels did they pick in all?
Answer by mananth(16949)

paul - 2 2/3 bushels = (3*2+2)=/3 = 8/3
..
Karl - 4 1/3 bushels = (3*4+1)/3 = 13/3
..
Carol - 3 2/3 bushels= (3*3 +2)/11 = 11/3
..
8/3 + 13/3 + 12/3 is the total
denominator is common =3
(8+13+12) /3= 33/2
= 16 1/2 bushels